OpenStudy (anonymous):

If i have 27 days to sell 15 apples, how many apples a day do i need to sell?

OpenStudy (anonymous):

well, you would do 27 divided by 15 and you would get 1.8 so yeah, about 2 :)

OpenStudy (anonymous):

course it does say 27 days and 15 apples so i am wrong

OpenStudy (amistre64):

at most, 15

OpenStudy (anonymous):

on average \(\frac{15}{27}=\frac{5}{9}\) a day

OpenStudy (amistre64):

lower the price ..

OpenStudy (anonymous):

this gotta be the most interesting problem of the day

OpenStudy (anonymous):

I think we are making this harder then it needs to be -_-

OpenStudy (amistre64):

bake a pie :) cut it into 27 pieces ....

OpenStudy (anonymous):

i wonder what answer the questioner is looking for? on average it is \(\frac{5}{9}\) but maybe you are supposed to say 1

OpenStudy (anonymous):

cant be 2 aday. 2 a day multiplied by 27 is 54..

OpenStudy (anonymous):

Ill posed question

OpenStudy (anonymous):

yeah i made a mistake. fixed it later though

OpenStudy (anonymous):

Why can't you just sell em all on day 1

OpenStudy (anonymous):

You would only like sell half a apple a day? This question doesn't make any sense